OpenStudy (anonymous):

What 3 consecutive even integers have a sum of -36?

OpenStudy (anonymous):

X + (X+2) + (X+4) = -36 solve for X X is our first even number, every consecutive will hence be X+2

OpenStudy (anonymous):

i dont understand the question

OpenStudy (godlovesme):

more help? :)

OpenStudy (anonymous):

YES PLEASE

OpenStudy (anonymous):

Consecutive means one after another. So if your first even number is 2, next will be 4, and next 6. Each is +2 the previous. So we select X as our first (as we dont know it's value)

OpenStudy (godlovesme):

@doctorwho123s consecutive integers r like 2, 4, 6 there is a 2 unit difference between each number

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so how do i solve the question ?

OpenStudy (godlovesme):

so if the first number is x the second should be x+2 and the third should be x+4

OpenStudy (godlovesme):

the sum is -36 so x+(x+2)+(x+4)=-36 can u solve for x @doctorwho123s

OpenStudy (anonymous):

you let your first number be X. Next will be X+2. Next will be X+2+2. Add them all up and equate to -36

OpenStudy (anonymous):

yes thank u, i understand know :)

OpenStudy (godlovesme):

welcome :) and nice job @ArsalanAijaz :)

OpenStudy (anonymous):

@Godlovesme how long till I get a title? :P

OpenStudy (godlovesme):

lol u can unlock "human calculator" :)

OpenStudy (anonymous):

how?

OpenStudy (godlovesme):

go to ur profile openstudy.com/users/arsalanaijaz

OpenStudy (anonymous):

let the numbers be x-2,x,x+2 be three consecutive integers ,then x-2+x+x+2=-36 3 x=-36 x=-12 so numbers are ?